OpenStudy (anonymous):

limit x to 0 of sin(4x)/2x

OpenStudy (anonymous):

Are you allowed to use L'Hopital's rule? If so, it's a very simple limit.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

\[\LARGE \lim_{x\to 0} \frac{\sin 4x}{2x} \] Multiply numerator and denominator by \[\LARGE \frac44\] you'll have.. \[\LARGE \lim_{x\to 0} \frac{4\cdot \sin 4x}{4\cdot 2x} \] \[\LARGE \lim_{x\to 0} \frac{\sin 4x}{4x} \cdot \frac{4}{2}\] You're supposed to know this rule: \[\LARGE \lim_{x\to 0}\frac{\sin x}{x}=1\] so the answer should be 2